# Defining implication relation for classical logic

### v11

@@ Abstract
In classical logic, “P implies Q” is equivalent to “not-P or Q”. It is well known that the equivalence is problematic. Actually, from “P implies Q”, “not-P or Q” can be inferred (“Implication-to-Disjunction” is valid), whereas from “not-P or Q”, “P implies Q” cannot be inferred in general (“Disjunction-to-Implication” is not generally valid), so the equivalence between them is invalid in general. This work aims to remove the incorrect Disjunction-to-Implication from classical logic (CL). The paper proposes a logic (IRL) with the expected properties: (a) CL is obtained by adding Disjunction-to-Implication to IRL, and (b) Disjunction-to-Implication is not derivable in IRL; while (c) fundamental laws in classical logic – including law of excluded middle and principle of double negation, law of non-contradiction and ex contradictione quodlibet, conjunction elimination and disjunction introduction, and hypothetical syllogism and disjunctive syllogism – are all retained in IRL.
Keywords
implication, conditional, material implication, material conditional, paradox, semantics, classical logic, propositional logic, first-order logic, proof system, Hilbert system, natural deduction
@@ 1 Introduction
In classical logic, “$P$ implies $Q$” ($P\to Q$) is taken to be logically equivalent to “it cannot be $P$ and not $Q$” ($\neg(P\wedge\neg Q)$) or “not-$P$ or $Q$” ($\neg P\vee Q$) by duality. This equivalence can be viewed as the definition of implication by disjunction and negation ($P\to Q=\neg P\vee Q$, called material implication), or used to define disjunction by implication and negation, so $P\to Q$ and $\neg P\vee Q$ can be replaced in a proof with one another.
@@
of classical logic means that any two propositions must be related by implication (total order). This does not agree with a large number of situations in the real world. For example, let $P=$ “Alice is in Athens” and $Q=$ “Alice is in London” – given an instant in the future, then both $P\to Q$ and $Q\to P$ cannot be true, so $(P\to Q)\vee(Q\to P)$ cannot be true.
Many efforts have been made to resolve this problem and important non-classical logics are developed, such as modal logics which introduce the “strict implication” and relevance logics which, as paraconsistent logics, reject ex contradictione quodlibet (ECQ).
@@
The paper explains why “Implication-to-Disjunction” ($(P\to Q)\to\neg P\vee Q$) is valid whereas “Disjunction-to-Implication” ($\neg P\vee Q\to(P\to Q)$) is not, by describing firstly the original meaning of “implication”. So the goal of this work is a logic in which Disjunction-to-Implication is removed while fundamental laws in classical logic such as law of excluded middle (LEM) and principle of double negation, law of non-contradiction (LNC) and ex contradictione quodlibet (ECQ), conjunction elimination and disjunction introduction, and hypothetical syllogism (transitivity of implication) and disjunctive syllogism, are all retained in IRL.
@@ 2.1 Four situations
Let $P$ and $Q$ be propositions.
Consider how $P$ affects $Q$. There are four possible situations when $P$ is true:
@@
“Coincidentally” means “not necessarily”, i.e. “without a certain mechanism”.
“Certain mechanism” means one of the two:
@@
    It is certain literally or physically (in some scope, cf. Subsection 4.6), see Example 4.6.2;
@@ 2.2 Original meaning of implication
According to this description, implication from propositions $P$ to $Q$ is (can be defined as) a relation determined by a certain mechanism (either physical or logical) from $P$ to $Q$, which guarantees that $Q$ is true whenever $P$ is true. Without such a mechanism, even if $Q$ is true when $P$ is true, $P\to Q$ is still false since $Q$ is just coincidentally true when $P$ is true.
@@ 2.3 Non-truth-functionality of implication
Let $L$ be the set of all the formulas representing a propositional language. Let $U$ be the set of all valuation functions from $L$ to the set of truth values $\{T,F\}$. Then:
According to the meaning of conjunction, there exists a function $f:\{T,F\}^{2}\to\{T,F\}$ such that for any $v\in U$ and any $\phi,\psi\in L$, $v(\phi\wedge\psi)=f(v(\phi),v(\psi))$, This is, $\wedge$ is truth-functional.
In contrast, according to the original meaning of implication, there exists no function $f:\{T,F\}^{2}\to\{T,F\}$ such that for any $v\in U$ and any $\phi,\psi\in L$, $v(\phi\to\psi)=f(v(\phi),v(\psi))$. This is, $\to$ is not truth-functional.
For instance, in terms of the meaning of conjunction, if we know that $P$ is true and $Q$ is true, then we know that $P\wedge Q$ is true for sure; whereas, in terms of the original meaning of implication, even if we know that $P$ is true and $Q$ is true, we are not sure about whether $P\to Q$ is true or false, because this depends on whether there is a “certain mechanism” from $P$ to $Q$ to guarantee that whenever $P$ is true $Q$ must be true.
From the semantic viewpoint, implication $\to$ is a binary relation rather than a binary operation like conjunction $\wedge$ and disjunction $\vee$. Specifically, the implication $\to$ being a binary relation on $L$, the truth value of $\phi\to\psi$ is determined by whether $(\phi,\psi)$ belongs to the set of the relation $\to$ that is a subset of $L\times L$ (semantically, it is the T-set of the implication, cf. Subsection 4.1), rather than by the truth value of $\phi$ and the truth value of $\psi$. We call $\to$ implication relation when emphasizing this fact.
@@
2.4 Incorrectness of material implication as “implication”
@@
$\neg P\vee Q$ is false” contains only one case: $P$ is true and $Q$ is false that includes two sub-cases, i.e. “$P$ is true and $Q$ is necessarily false” and “$P$ is true and $Q$ is coincidentally false”. The two sub-cases is a subset of situations where $P\to Q$ is false. In symbols, ¬($\neg(\neg P\vee Q)\to\neg(P\to Q)$ holds whereas $\neg(P\to Q)\to\neg(\neg P\vee Q)$ does not. This result is the same as above in terms of contraposition.
@@ 2.5 Difference from “strict implication”
The negation of strict implication, “it is possible not that $P$ implies $Q$”, is not excluding the situation “it is possible that $P$ implies $Q$”. This is different from the original meaning of implication defined previously.
In fact, “strict implication” does not define the meaning of the implication itself: let $S=P\to Q$, the strict implication $\Box(P\to Q)=\Box S$ says just “necessarily $S$”, nothing about the implication $P\to Q$ itself.
@@ 2.6 Construction of the right logic
    Classical scope. The system is “in the classical scope”. For example, it is a two-valued logic and a sub-system of classical logic so that any operators new to classical logic like modal operators are not considered.
@@
    Strength. Under the above two conditions, the system should be as strong as possible. For example, both LEM and ECQ are not rejected. LEM is irrelevant to the original meaning of implication. ECQ, which is necessary for retaining both disjunction introduction and disjunctive syllogism (cf. Section 6), does not conflict with the original meaning of implication (cf. Subsection 4.1). And Implication-to-Disjunction $(\phi\to\psi)\to\neg\phi\vee\psi$ is kept since it follows directly from the original meaning of implication.
There are many choices of ways for constructing the system. Two main approaches are:
@@
Because of its symmetry, bi-implication-based approach has an advantage in structure. The logic IRL presented in Section 3 and its algebraic system IRLA in Section 4, are constructed in the bi-implication-based approach.
@@ 3.1 Definitions and notations
is defined as the union of $X\Rightarrow Y$ and $Y\Rightarrow X$ (view a single-conclusion rule as a singleton).
@@
Meta rule: If a premise or a conclusion of a rule contains symbol(s) not belonging to the object language, then the rule is a meta rule, e.g. the rule of conditional proof $(\phi\vdash\psi)\Rightarrow(\vdash\phi\to\psi)$ is a meta rule.
Domain-specific rule: A domain-specific rule is sometimes called an axiom or a theorem in that domain. For example, in the Boolean lattice, $x\leq y\Leftrightarrow x\wedge y=x$ is listed as an axiom although it is essentially a (bi-inference) domain-specific rule.
@@
A proof of $y$ from $X$, denoted $X\vdash y$, is a non-empty finite sequence of formulas $(y_{1},...,y_{n})$ with $y_{n}=y$, where $y_{k}\;(1\leq k\leq n)$ is an axiom, a theorem, or a result from a subset of $X\cup\{y_{1},...,y_{k-1}\}$ by applying a rule of inference. The length or steps of a proof is the length of the proof sequence.
An unconditional proof is a proof with an empty set of premises. Formula $y$ is called a theorem if there exists an unconditional proof of $y$, denoted $\vdash y$.
A Conditional proof is a proof with a non-empty set of premises.
@@
All axioms are theorems.
In a proof sequence $(y_{1},...,y_{n})$ of an unconditional proof, each subsequence of the form $(y_{1},...,y_{k})\,(1\leq k\leq n)$ is a unconditional proof, so that every member $y_{k}\,(1\leq k\leq n)$ is a theorem.
@@
Let $X$ and $Y$ be non-empty finite sets of formulas.
A rule $X\Rightarrow Y$ is derivable if $X\vdash Y$. This can be expressed by the meta rule $(X\vdash Y)\Rightarrow(X\Rightarrow Y)$.
Obviously, any primitive rule is derivable.
@@
Let $x_{1},...,x_{k}$ be variables, $t_{1},...,t_{k}$ be terms, and $E$ be an expression.
@@
The rule of uniform substitution is of the form $E\Rightarrow E\{t_{1}/x_{1},...,t_{k}/x_{k}\}$, where $E$ is an axiom, a theorem, a rule, or a meta rule of a formal system.
Remark 3.1.3.
@@
The presentation format of proof is as follows.
A unconditional proof $\vdash\psi$ of length $n$ is presented in the format:
@@
$m$ $\vdash\phi_{m}$ Premise
$m+1$ $\vdash\psi_{1}$ reason
@@
$m+n-1$ $\vdash\psi_{n-1}$ reason
$m+n$ $\vdash\psi$ reason
Note: In some expressions, “$\alpha_{0}\backsim\alpha_{1}\backsim\cdots\backsim\alpha_{n}$” is a shorthand of “$\alpha_{0}\backsim\alpha_{1}$, $\alpha_{1}\backsim\alpha_{2}$, …, $\alpha_{n-1}\backsim\alpha_{n}$”, so that in a proof the format
$\alpha_{0}$ $1$     $\backsim\alpha_{1}$vdots     $\vdots$ $n$     $\backsim\alpha_{n}$
@@
$1$ $\alpha_{0}\backsim\alpha_{1}$vdots $\vdots$ $n$ $\alpha_{n-1}\backsim\alpha_{n}$
[…]
A note on nested proof
By definition of proof (Definition 3.1.2), any member of a proof sequence must be either a theorem (or an assumed theorem in its outer layer) or a result from premise(s) in the same layer by a rule (or by an assumed rule in its outer layer), not allowing any other cases. Attention must be paid to this, especially in nested conditional proofs.
@@
Conditional proof of $\phi\to(\psi\to\phi)$:
1     $\phi$ Premise (main) 2       $\psi$ Premise (nested) 3       $\phi$ 1: Reiteration – Not allowed 4     $\psi\to\phi$ 2–3: Rule of conditional proof 5 $\phi\to(\psi\to\phi)$ 1–4: Rule of conditional proof
Line 3 is not allowed because $\phi$ in this line does not belong to any of the allowed cases: (a) it is theorem, or it is assumed be a theorem in its outer premises; (b) it is a result from premises of its own layer (that must be $\psi$ in line 2) by applying a rule.
@@
Conditional proof of $(\top\to\phi)\to(\psi\to\phi)$:
1     $\top\to\phi$ Premise (main) 2       $\psi$ Premise (nested) 3       $\top$ A theorem ($\top$ is a theorem) 4     $\psi\to\top$ 2–3: Rule of conditional proof 5     $\psi\to\phi$ 4,1: Rule of transitivity 6 $(\top\to\phi)\to(\psi\to\phi)$ 1–5: Rule of conditional proof
The following details a meta proof of a meta rule.

### v12

@@ Abstract
In classical logic, “P implies Q” is equivalent to “not-P or Q”. It is well known that the equivalence is problematic. Actually, from “P implies Q”, “not-P or Q” can be inferred (“Implication-to-Disjunction” is valid), whereas from “not-P or Q”, “P implies Q” cannot be inferred in general (“Disjunction-to-Implication” is not generally valid), so the equivalence between them is invalid in general. This work aims to remove the incorrect Disjunction-to-Implication from classical logic (CL). The logical system (the logic IRL) this paper proposes has the expected properties: (a) CL is obtained by adding Disjunction-to-Implication to IRL, and (b) Disjunction-to-Implication is not derivable in IRL; while (c) fundamental laws in classical logic, including law of excluded middle (LEM) and principle of double negation, law of non-contradiction (LNC) and ex contradictione quodlibet (ECQ), conjunction elimination and disjunction introduction, and hypothetical syllogism and disjunctive syllogism, are all retained in IRL.
Keywords implication, conditional, material implication, paradox, classical logic, propositional logic, first-order logic, proof system, Hilbert system, natural deduction, semantics, algebraic logic
@@ 1 Introduction
In classical logic, “$P$ implies $Q$” ($P\to Q$) is taken to be logically equivalent to “it cannot be $P$ and not $Q$” ($\neg(P\wedge\neg Q)$) or “not-$P$ or $Q$” ($\neg P\vee Q$) by duality. This equivalence can be viewed as the definition of implication by disjunction and negation ($P\to Q=\neg P\vee Q$, called material implication), or used to define disjunction by implication and negation, so $P\to Q$ and $\neg P\vee Q$ can be replaced with one another in a derivation.
@@
of classical logic means that any two propositions must be related by implication (total order). This does not agree with a large number of situations in the real world. For example, let $P=$ “Alice is in Athens” and $Q=$ “Alice is in London” (given an instant in the future), then both $P\to Q$ and $Q\to P$ cannot be true, so $(P\to Q)\vee(Q\to P)$ cannot be true.
Many efforts have been made to resolve this problem and important non-classical logics are developed, such as modal logics which introduce the “strict implication”, relevance logics which require that the antecedent and consequent of implications to be relevantly related (so they are paraconsistent logics which reject ex contradictione quodlibet (ECQ)).
@@
The paper explains why “Implication-to-Disjunction” ($(P\to Q)\to\neg P\vee Q$) is valid whereas “Disjunction-to-Implication” ($\neg P\vee Q\to(P\to Q)$) is not, by describing firstly the original meaning of “implication”. So the goal of this work is a logic in which Disjunction-to-Implication is removed while fundamental laws in classical logic such as law of excluded middle (LEM) and principle of double negation, law of non-contradiction (LNC) and ex contradictione quodlibet (ECQ), conjunction elimination and disjunction introduction, and hypothetical syllogism (transitivity of implication) and disjunctive syllogism, are all retained.
@@ 2.1 Four situations
Let $P$ and $Q$ be propositions. Consider how $P$ affects $Q$. There are four possible situations about $Q$ when $P$ is true:
@@
  • “coincidentally” means “without a certain mechanism”.
And “certain mechanism” means one of the two:
@@
    It is certain literally or physically (in some scope, cf. Subsection 4.6), for example, “in the open air” is a physical mechanism of “if it rains, then the ground gets wet” (cf. Example 4.6.2);
@@ 2.2 Original meaning of implication
According to this description, implication from propositions $P$ to $Q$ is (can be defined as) a relation determined by a certain mechanism (physical or logical) from $P$ to $Q$, which guarantees that $Q$ is true whenever $P$ is true. Without such a mechanism, even if $Q$ is true when $P$ is true, $P\to Q$ is still false since $Q$ is just coincidentally true when $P$ is true.
@@ 2.3 Non-truth-functionality of implication
Let $L$ be the set of all the formulas representing a propositional language. Let $U$ be the set of all valuation functions from $L$ to the set of truth values $\{\mathrm{T,F}\}$. Then:
According to the meaning of conjunction, there exists a function $f:\{\mathrm{T,F}\}^{2}\to\{\mathrm{T,F}\}$ such that for any $v\in U$ and any $\phi,\psi\in L$, $v(\phi\wedge\psi)=f(v(\phi),v(\psi))$, That is, $\wedge$ is truth-functional.
In contrast, according to the original meaning of implication, there exists no function $f:\{\mathrm{T,F}\}^{2}\to\{\mathrm{T,F}\}$ such that for any $v\in U$ and any $\phi,\psi\in L$, $v(\phi\to\psi)=f(v(\phi),v(\psi))$. That is, $\to$ is not truth-functional.
For instance, in terms of the meaning of conjunction, if we know that $P$ is true and $Q$ is true, then we know for sure that $P\wedge Q$ is true; whereas, in terms of the original meaning of implication, even if we know that $P$ is true and $Q$ is true, we are not sure about whether $P\to Q$ is true or false, because this depends on whether there is “a certain mechanism” from $P$ to $Q$ to guarantee that whenever $P$ is true $Q$ must be true.
In fact, we know surely the truth value of $P\wedge Q$ for all possible four combinations of truth values $\{\mathrm{TT,TF,FT,FF}\}$ of $P$ and $Q$, whereas we do not know the truth value of $P\to Q$ from that of $P$ and $Q$ except for the only one combination $\mathrm{TF}$: only when $P$ is true and $Q$ is false, we know for sure that $P\to Q$ is false, because (only) $P\wedge\neg Q\to\neg(P\to Q)$ is valid according to the original meaning of implication (cf. Subsection 2.4).
From the semantic viewpoint, implication $\to$ is a binary relation rather than a binary operation like conjunction $\wedge$ and disjunction $\vee$. Specifically, the implication $\to$, as a binary relation on $L$, the truth value of $\phi\to\psi$ is determined by whether $(\phi,\psi)$ belongs to the set of the relation $\to$ that is a subset of $L\times L$ (semantically, it is just the T-set of the implication, cf. Subsection 4.1), rather than by the truth value of $\phi$ and the truth value of $\psi$. We call $\to$ implication relation when emphasizing this fact.
@@
2.4 Incorrectness of “material implication” as implication
@@
$\neg P\vee Q$ is false” contains only one case: $P$ is true and $Q$ is false that includes two sub-cases, i.e. “$P$ is true and $Q$ is necessarily false” and “$P$ is true and $Q$ is coincidentally false”. The two sub-cases is a subset of situations where $P\to Q$ is false (cf. Subsection 2.2). In symbols, ¬($\neg(\neg P\vee Q)\to\neg(P\to Q)$ (i.e. $P\wedge\neg Q\to\neg(P\to Q)$) holds, whereas $\neg(P\to Q)\to\neg(\neg P\vee Q)$ does not. This result is the same as above in terms of contraposition.
@@ 2.5 Difference from “strict implication”
The negation of strict implication, “it is possible not that $P$ implies $Q$”, is not excluding the situation “it is possible that $P$ implies $Q$”. This is different from the original meaning of implication defined in previous Subsection 2.2.
As a matter of fact, “strict implication” does not define the meaning of implication itself: let $S=P\to Q$, the strict implication $\Box(P\to Q)=\Box S$ says just “necessarily $S$”, nothing about the implication $P\to Q$ itself.
@@ 2.6 Construction of the right logic
    Classical scope. The logical system to be constructed is “in the classical scope”. For example, it is a two-valued logic and a sub-system of classical logic, so that any operators new to classical logic like modal operators are not considered.
@@
    Strength. Under the above two conditions, the system should be as strong as possible. For example, both LEM and ECQ are not rejected. LEM is irrelevant to the original meaning of implication. ECQ, which is necessary for retaining both disjunction introduction and disjunctive syllogism (cf. Section 9), does not conflict with the original meaning of implication (cf. Subsection 4.1). And Implication-to-Disjunction $(\phi\to\psi)\to\neg\phi\vee\psi$ is kept since it follows directly from the original meaning of implication as explained previously.
There are many choices of ways for constructing logics. Two main approaches are:
@@
Because of its symmetry, bi-implication-based approach has an advantage in structure. The logic IRL presented in Section 3 and its algebraic system IRLA in Section 4, are constructed in the bi-implication-based approach. The natural deduction for IRL is constructed in the implication-based approach that is more suitable to it.
@@ 3.1 Definitions and notations
is defined as the union of $X\Rightarrow Y$ and $Y\Rightarrow X$ (treat a single-conclusion rule as a singleton for the union operation).
@@
Domain-specific rule: A domain-specific rule is sometimes called an axiom or a theorem in that domain. For example, in a Boolean lattice extended from a Boolean algebra, $x\leq y\Leftrightarrow x\wedge y=x$ is usually listed as the defining axiom for its partial order ($\leq$), although it is essentially a (bi-inference) domain-specific rule.
@@
A proof of $y$ from $X$, denoted
$X\vdash y,$
is a non-empty finite sequence of formulas $(y_{1},...,y_{n})$ with $y_{n}=y$, where $y_{k}\;(1\leq k\leq n)$ is an axiom, a theorem, or a result from a subset of $X\cup\{y_{1},...,y_{k-1}\}$ by applying a rule of inference.
  • An unconditional proof is a proof with an empty set of premises. Formula $y$ is called a theorem if there exists an unconditional proof of $y$, denoted
    $\vdash y.$
  • A Conditional proof is a proof with a non-empty set of premises, i.e. $X\vdash y$ where $X$ is non-empty. Note that a member of the set $X$ of premises is not a member of its proof sequence $(y_{1},...,y_{n})$.
The length or steps of a proof is the length of its proof sequence.
@@
  • All axioms are theorems. Therefore, any member of a proof sequence is just either a theorem or a result from a subset of premises and previous members of the proof sequence by applying a rule.
  • In a proof sequence $(y_{1},...,y_{n})$ of an unconditional proof, each subsequence of the form $(y_{1},...,y_{k})\,(1\leq k\leq n)$ is a unconditional proof, thus every member $y_{k}\,(1\leq k\leq n)$ is a theorem.
@@
Let $X$ and $Y$ be finite sets of formulas where $Y$ is non-empty.
A rule $X\Rightarrow Y$ is derivable if and only if there exists a proof $X\vdash Y$ where rules used are only existing ones. An existing rule is a primitive rule or a derived rule.
In terms of the definitions of proof and derivable rule, any primitive rule is derivable.
Remark 3.1.3 (Meta rule and meta proof).
If a premise or a conclusion of a rule or a proof contains a symbol not belonging to the object language, then the rule or the proof is a meta rule or meta proof. For example, the rule $(\phi\vdash\psi)\Rightarrow(\vdash\phi\to\psi)$ of conditional proof is a meta rule, a proof $(\phi\vdash\psi)\vdash(\vdash\phi\to\psi)$ of the deduction theorem is a meta proof.
@@
Let $x_{1},...,x_{k}$ be variables, $t_{1},...,t_{k}$ be terms, and $E$ be an expression containing variables.
@@
The rule of uniform substitution is the form $E\Rightarrow E\{t_{1}/x_{1},...,t_{k}/x_{k}\}$, where $E$ is an axiom, a theorem, a rule, or a meta rule of a formal system.
@@
The presentation format of a proof is as follows.
An unconditional proof $\vdash\psi$ of length $n$ is presented in the format:
@@
$m$ $\phi_{m}$ Premise (Assumption)
$m+1$ $\psi_{1}$ reason
@@
$m+n-1$ $\psi_{n-1}$ reason
$m+n$ $\psi$ reason
A proof (unconditional or conditional) can has other proofs (unconditional or conditional) embedded within it. In the presentations, this nested structure of proofs must be clearly indicated in a consistent manner such as indentation used in this paper.
An example that an unconditional proof has an inner conditional proof:
1 $\psi_{11}$ theorem-name
2     $\phi_{21}$ Premise
3     $\psi_{21}$ theorem-name
4     $\psi_{22}$ 2,3: rule-name (From $\{2,3\}$ by rule-name)
5 $\phi_{21}\to\psi_{22}$ 2–4: Rule of conditional proof
6 $\psi_{12}$ 1,5: rule-name
Note that the result of a conditional proof (by the rule of conditional proof) is not a member of its own proof sequence, so it must be unindented to just one level higher (one layer outer) as shown in the example (line 5).
[…]
Hence, in the presentation of a proof, the format
$\alpha_{0}$
$1$     $\backsim\alpha_{1}$
$\vdots$     $\vdots$
$n$     $\backsim\alpha_{n}$
@@ 6 $\psi_{12}$ 1,5: rule-name
$1$ $\alpha_{0}\backsim\alpha_{1}$
$\vdots$ $\vdots$
$n$ $\alpha_{n-1}\backsim\alpha_{n}$
A note on nested proofs
By definition (Definition 3.1.2, and cf. Remark 3.1.2), any member of a proof sequence must be either a theorem or a result from a subset of premises and previous members by a rule. Specifically, let $X\vdash y_{n}$ has a proof sequence $(y_{1},...,y_{n})$, then any $y_{k}$ must be in one of the following two cases:
  • $y_{k}$ is a theorem, or it is assumed be a theorem in a premise of an outer proof, or it is a theorem derived in an outer proof;
  • $y_{k}$ is obtained from a subset of $X\cup\{y_{1},...,y_{k-1}\}$ by applying a rule, or by an assumed rule in a premise of an outer proof, or by a rule derived in an outer proof.
No other cases are allowed.
This is further illustrated in Table 1.
Attention must be paid to this, especially for a nested conditional proof.
Table 1: Only two cases allowed for members of proof sequence
An outer proof The inner proof Comment
$A_{1}$
$\vdots$
$A_{m}$$\{A_{1},...,A_{m}\}$ is the set of premises/assumptions of the outer proof a
$B_{1}$
$\vdots$
$B_{i}$ For $B_{i}$, a member of the outer proof sequence, only two cases are allowed: it is a theorem, or it is a result from a subset of $\{A_{1},...,A_{m}\}\cup\{B_{1},...,B_{i-1}\}$ by applying a rule b
$\vdots$
$B_{j}$
$P_{1}$
$\vdots$
$P_{n}$$\{P_{1},...,P_{n}\}$ is the set of premises/assumptions of the inner proof a
$Q_{1}$
$\vdots$
$Q_{k}$ For $Q_{k}$, a member of the inner proof sequence, only two cases are allowed: (a) it is a theorem, or it is assumed be a theorem in $\{A_{1},...,A_{m}\}$, or it is a theorem in $\{B_{1},...,B_{j}\}$; or (b) it is a result from a subset of $\{P_{1},...,P_{n}\}\cup\{Q_{1},...,Q_{k-1}\}$ by a rule, or by a rule assumed in $\{A_{1},...,A_{m}\}$, or by a rule in $\{B_{1},...,B_{j}\}$
$\vdots$
  • a
    If the set of premises is empty, it is an unconditional proof; otherwise, it is a conditional proof.
  • b
    If the outer proof itself is embedded in another proof, treat it the same way as the inner proof in the table.
@@
Proof of $\phi\to(\psi\to\phi)$:
1     $\phi$ Premise (outer) 2       $\psi$ Premise (inner) 3       $\phi$ 1: Reiteration – Not allowed 4     $\psi\to\phi$ 2–3: Rule of conditional proof 5 $\phi\to(\psi\to\phi)$ 1–4: Rule of conditional proof
Line 3 is not allowed because $\phi$ in this line does not belong to any of the allowed cases: (a) it is a theorem or a theorem indicated in an outer proof ($\phi$ is not assumed be a theorem in the outer premise i.e. line 1); (b) it is a result from a subset of premises and previous members of its own layer (that must be a subset of $\{\psi\}$ from line 2) by applying a rule or by an rule in an outer proof.
@@
Proof of $(\top\to\phi)\to(\psi\to\phi)$:
1     $\top\to\phi$ Premise (outer) 2       $\psi$ Premise (inner) 3       $\top$ A theorem ($\top$ is indeed a theorem) 4     $\psi\to\top$ 2–3: Rule of conditional proof 5     $\psi\to\phi$ 4,1: Rule of transitivity 6 $(\top\to\phi)\to(\psi\to\phi)$ 1–5: Rule of conditional proof
Sometimes, a meta proof (for a meta language) contains an ordinary proof (for an object language) such as the following example, which details a meta proof of a meta rule.